40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Feb 9 18:33:59 2010, Frank, Laser, RC noise, even newer results refcav-noise_2010-02-09.png
    Reply  Wed Feb 10 04:01:50 2010, rana, Laser, RC noise, even newer results 
Message ID: 62     Entry time: Tue Feb 9 18:33:59 2010     Reply to this: 63
Author: Frank 
Type: Laser 
Category: RC noise 
Subject: even newer results 

measured a couple of times today with everything re-aligned and different gain settings for the FSS stuff. Measured also to lower frequencies. The problem here is that the frequency band of interest can not be measured with epics (to slow) and the measurement using the SR785 takes so long that the operating point of the VCO and therefore the coefficient changes during the measurement and different spectra don't fit together quite well. so we have to measure a couple of times to get some measurements fitting together when the changes are so small that we don't see it with our eyes. this gives us a new upper limit of:


maximum gain settings:
CG : 24dB
FG : 13dB

ELOG V3.1.3-